As an investor, you decided to purchase a-90 day Bank Bill that has a face value of $100,000. When purchased, the bill had 68 days to maturity and was yielding 8.25%. After holding the bill for 10 days, interest rates increased to 9.00% and you sold the bill. What dollar amount of interest did you earn?

  1. -$ 11.25
  2. $ 25.78
  3. $ 87.65
  4. $103.75
  5. $109.89
Answer #1

Answer:

Correct answer is:

d. $103.75

Explanation:

When purchased:

Time to maturity = 68 days

Yield = 8.25%

Price purchased = $100,000 / (1 + 8.25% * 68/365) = $98,486.28

Sale of bill:

Sale of Bill after holding 10 days:

Time to maturity = 68 - 10 = 58 days

Yield = 9.00%

Sale value = $100,000 / (1 + 9% * 58/365) = $98,590.03

Interest earned = $98,590.03 - $98,486.28 = $103.75

As such option d is correct and other options a, b, c and e are incorrect.